# Multi-class Temporal Logic Neural Networks

### v1

@@ IV-C2 Loss-based distance
Since the margin is introduced only to improve the classifier during the training process, the loss-based distance is the loss function while removing the margin. As a result, the notion of the margin is not used in the prediction.
1:Time-series dataset $\mathcal{S}$, candidate sub-formulae $\varphi$, multi-conjunction-disjunction matrix $M_{m}$, ECOC coding matrix $E$, number of iterations $K$.
@@
3:Construct a multi-class NN-TLI based on the candidate sub-formulae $\varphi$ and multi-conjunction-disjunction matrix $M_{m}$. Initialize the parameters.
@@ V Case study
In this section, we first implement our approach to classify and interpret multiple behaviors in a naval scenario using STL formulae and show how we can learn interpretable attributes of each class using the ECOC coding matrix. We also compare with the state-of-the-art methods on a synthetic dataset and demonstrate the benefits of our method. For all the case studies in this section, we test both the attribute-based classification (Problem 1) and the class-based classification methods (Problem 2). The mean and variance of training time and misclassification rate are similar for both methods, which implies that there is no fundamental difference between these two methods. Due to the space limitation, we only report results from the attribute-based classification in this section.
@@ V-A Naval Scenario Dataset
In the naval scenario, the dataset consists of trajectories of three classes. The trajectories are two-dimensional and have $60$ time steps in total. As shown in Fig 2, the classes can be described as class 1: “arrive harbor while not reaching island”; class $2$: “arrive harbor while passing through island”; class $3$: “return to the open sea without reaching harbor or island.” Class $1$ has $1000$ trajectories, class $2$ and class $3$ have $500$ trajectories each.
In this experiment, we use two kinds of coding matrices. The first kind of matrix (Table IIa) is coded based on the class. In this case, traditional class learning is applied, but not attribute learning; each binary classifier $f$ is used to classify one class. We do not learn the attributes but only the property of each class. The second kind of matrix (Table IIIa,IVa,Va) is coded based on the attributes. Each binary classifier $f$ is used to classify one attribute. The trajectories are classified based on whether they contain certain attributes. For example, in $E_{a1}$, class $1$ has attribute $1$ but not attribute $2$; class $2$ has both attribute $1$ and $2$; class $3$ has neither attribute $1$ or $2$. According to [17], there are at most $12^{c-1}-$ 1 usable columns for a $c$-class problem after excluding the complementary, all-zeros or all-ones columns. We used different attribute-based coding matrices $E_{a1}$, $E_{a2}$ and $E_{a3}$ to show that we can achieve semantically equivalent results regardless of the design of the coding matrix.
@@
where $N$ is the number of samples in the dataset. The results are shown in Table II-V. Using the attribute-based coding matrix, we can reach zero misclassification rate with less training time since there are fewer binary classifiers to learn compared to the class-based coding matrix. The attribute learning can also provide more concise STL formulae to interpret the features that are easy for humans to understand.
@@
$f_{1}$
$f_{2}$
$f_{3}$
Class $1$
Class $2$
Class $3$
@@
Class
STL formulae
Description
Class 1
$\lozenge_{[4,60]}(x<25.0)\land\square_{[5,28]}(y>23.6)$
reach harbor and not reach island
Class 2
$\lozenge_{[0,26]}(y<21.0)$
reach island
Class 3
$\square_{[0,59]}(x>34.6)$
stay in open sea
@@
Attribute 1 $\lozenge_{[3,60]}(x<24.3)$ reach harbor
Attribute 2 $\lozenge_{[0,50]}(y<22.3)$ reach island
@@
Attribute
STL formulae
Description
Attribute 1
$\lozenge_{[0,60]}(x<23.1)$
reach harbor
Attribute 2
$\lozenge_{[0,60]}(x<25.3)\land\square_{[12,42]}(y>23.3)$
reach harbor and not reach island
@@
Attribute
STL formulae
Description
Attribute 1
$\square_{[10,53]}(y>21.9)$
not reach island
Attribute 2
$\square_{[9,51]}(y>22.5)\land\lozenge_{[17,60]}(x<26.9)$
not reach island and reach harbor
@@ V-B Synthetic Dataset
$p=(\alpha\leq x\leq\beta)\land\gamma\leq y\leq\delta),\$ (20)
@@
$\phi_{4}$ $=\lozenge_{[10,40]}(-5\leq x\leq-3\land-5\leq y\leq-3),$
Figure 3 shows some example trajectories in the synthetic dataset. The signal is required to arrive at some regions within the desired time interval specified by the STL specifications corresponding to the class.
The synthetic dataset contains 5 classes labeled as $c_{1}$, $c_{2}$, $c_{3}$, $c_{4}$, $c_{5}$ with trajectories individually satisfy:
@@
fold
$mSTL$
$dt\Delta$
$rnn$
$mSTL$
$dt\Delta$
$rnn$
$mSTL$
$dt\Delta$
$rnn$
$mSTL$
$dt\Delta$
$rnn$
$mSTL$
$dt\Delta$
$rnn$
$mSTL$
$dt\Delta$
$rnn$
1
0.000
0.000
0.000
32
324
7
0.000
0.000
0.000
38
1689
8
0.000
0.000
0.000
58
3236
11
2
0.000
0.000
0.000
38
299
6
0.000
0.000
0.000
45
1431
8
0.000
0.000
0.000
123
3220
11
3
0.000
0.000
0.000
36
427
7
0.000
0.000
0.000
67
1786
8
0.000
0.000
0.000
73
3360
11
4
0.000
0.000
0.000
86
395
6
0.000
0.000
0.000
34
1594
8
0.000
0.000
0.000
86
3821
10
5
0.000
0.000
0.000
42
359
17
0.000
0.000
0.000
89
1931
18
0.000
0.000
0.000
61
3353
20
0.000
0.000
0.000
47
361
22
0.000
0.000
0.000
55
1686
10
0.000
0.000
0.000
80
3398
13
@@
We conduct experiments on the datasets with sample sizes of $500$, $2500$ and $5000$; these three datasets have $100$, $500$ and $1000$ samples for each class, respectively. We also compare the performance with the decision-tree-based algorithm [21] and a recurrent neural network [22] on these datasets. The decision-tree-based method ($dt\Delta$) uses a decision tree with a maximum depth of 5. The recurrent neural network (denoted as $rnn$ in Table VI) has $3$ hidden layers, with the first layer as an LSTM layer.
Table VI shows the results of these three methods on a 5-fold cross-validation. The ECOC matrix we use is the class-based coding matrix since $dt\Delta$ algorithm cannot implement the attribute learning. The results show that the multi-class NN-TLI, called $mSTL$ for short reference, can learn the STL formulae describing the properties of each class and achieve a low misclassification rate with high efficiency. The $mSTL$ and $rnn$ method can achieve much higher efficiency than the $dt\Delta$, while $rnn$ cannot provide the characteristics of trajectories of each class. The decision tree-based algorithm $dt\Delta$ is able to learn the STL formulae. However, the training time of $dt\Delta$ increases linearly with the sample size of the dataset. The reason for that is the computational bottleneck of the decision tree method, as it iterates all the input instances to find the optimal split at a given node. This step is repeated for each node of the tree [23]. Thus, the computational time grows linearly with the total number of instances. However, the neural network method iterates over the selected batch data at each epoch, which is small relative to the total number of instances.
@@
To investigate the effectiveness of the margin in the loss function during the training process, we recorded the average absolute robustness value of the correctly classified trajectories shown in Fig 4. For comparison, we also trained a multi-class NN-TLI without margin in the loss function and recorded the average absolute robustness, also shown in Fig 4. It can be observed that the introduction of margin to the loss can increase the absolute value of the average robustness. In other words, the margin can improve the separation between the trajectories during the training process and achieve better performance.
We also implement a zero-shot learning [24] to show the power of attribute learning using multi-class NN-TLI. The multi-class NN-TLI is trained over the observed class $c_{1}$, $c_{2}$, $c_{3}$, $c_{4}$ using the coding matrix $E_{obs}$ in Table VIIa and learn $4$ binary classifier $f_{1}$, $f_{2}$, $f_{3}$, $f_{4}$ correlating to $4$ kinds of attributes. These four attributes are interpreted by four STL formulae:
@@
$\psi_{4}$ $=\lozenge_{[1,20]}(-4.1<x<-0.2\land-3.8<y<-0.5)$
we then use these attributes to predict the unobserved class $c_{5}$ labeled as $E_{pred}$ in Table VIIb. We can achieve the $0.00$ $MCR$ for classifying the trajectories in $c_{5}$. We can claim that the multi-class NN-TLI is able to capture the attributes of the signals and interpret the characteristics by the learned STL formulae.
$f_{1}$
$f_{2}$
$f_{3}$
$f_{4}$
$c_{1}$
$c_{2}$
$c_{3}$
$c_{4}$
(a) The coding matrix $E_{obs}$
$f_{1}$
$f_{2}$
$f_{3}$
$f_{4}$
$c_{5}$
(b) The coding matrix $E_{pred}$
TABLE VII: The ECOC coding matrix for zero-shot learning.
@@ VI Conclusions and Future Work
We propose a neural network for multi-class classification problems on time-series data. The network can provide STL formulae to interpret the attributes of signals and classify the class of signals based on the attribute representation. In the future, we will focus on determining the importance of each attribute by learning the weight of each attribute in multi-class problems.

### v2

@@ IV-C2 Loss-based distance
Since the margin is introduced only to improve the classifier during training, the loss-based distance is the loss function while removing the margin. As a result, the notion of the margin is not used in the prediction.
1:Time-series dataset $\mathcal{S}$, ECOC coding matrix $E$, number of iterations $K$.
@@
3:Initialize the parameters of multi-class NN-TLI.
@@ V Case study
In this section, we first implement our approach to classify and interpret multiple behaviors in a naval scenario using STL formulae and show how we can learn interpretable attributes of each class using the ECOC coding matrix. We also compare with state-of-the-art methods on a synthetic dataset and demonstrate the benefits of our method. For all the case studies in this section, we test both the attribute-based classification (Problem 1) and the class-based classification methods (Problem 2). The mean and variance of training time and misclassification rate are similar for both methods, which implies that there is no fundamental difference between these two methods. Due to the space limitation, we only report results from the attribute-based classification in this section.
@@ V-A Naval Scenario Dataset
We use a dataset from [3] related to a naval surveillance scenario. The dataset consists of three kinds of vessel behaviors. The trajectories are two-dimensional and have $60$ time steps in total. As shown in Fig 2, the trajectories are divided into $3$ classes described as class 1: “arrive harbor while not reaching island”; class $2$: “arrive harbor while passing through island”; class $3$: “return to the open sea without reaching harbor or island.” Class $1$ has $1000$ trajectories, class $2$ and class $3$ have $500$ trajectories each.
In this experiment, we use two kinds of coding matrices. The first kind of matrix (Table IIa) is coded based on the class. In this case, traditional class learning is applied, but not attribute learning; each binary classifier $f$ is used to classify one class. We do not learn the attributes but only the property of each class. The second kind of matrix (Table IIIa,IVa,Va) is coded based on the attributes. Each binary classifier $f$ is used to classify one attribute. The trajectories are classified based on whether they contain certain attributes. For example, in $E_{a1}$, class $1$ has attribute $1$ but not attribute $2$; class $2$ has both attribute $1$ and $2$; class $3$ has neither attribute $1$ or $2$. According to [19], there are at most $12^{c-1}-$ 1 usable columns for a $c$-class problem after excluding the complementary, all-zeros or all-ones columns. We used different attribute-based coding matrices $E_{a1}$, $E_{a2}$, and $E_{a3}$ to show that we can achieve semantically equivalent results regardless of the design of the coding matrix.
@@
where $N$ is the number of samples in the dataset. The initialization of the multi-class NN-TLI, the batch size, and the number of training epochs are consistent for all cases. The results are shown in Table II-V. Using the attribute-based coding matrix, we can reach zero misclassification rate with less training time since there are fewer binary classifiers to learn compared to the class-based coding matrix. The attribute learning can also provide more concise STL formulae to interpret the features that are easy for humans to understand.
@@
$f_{1}$ $f_{2}$ $f_{3}$
Class $1$ + - -
Class $2$ + -
Class $3$ +
@@
Class STL formulae Description
Class 1 $\lozenge_{[2,60]}(x<22.5)\land\square_{[7,45]}(y>23.4)$ reach harbor and not reach island
Class 2 $\lozenge_{[10,49]}(y<20.0)$ reach island
Class 3 $\square_{[7,43]}(x>33.8)$ stay in open sea
@@
Attribute 1 $\lozenge_{[7,60]}(x<22.1)$ reach harbor
Attribute 2 $\lozenge_{[0,48]}(y<19.8)$ reach island
@@
Attribute STL formulae Description
Attribute 1 $\lozenge_{[6,60]}(x<23.1)$ reach harbor
Attribute 2 $\lozenge_{[4,60]}(x<22.3)\land\square_{[8,44]}(y>23.4)$ reach harbor and not reach island
@@
Attribute STL formulae Description
Attribute 1 $\square_{[9,55]}(y>22.4)$ not reach island
Attribute 2 $\square_{[4,45]}(y>23.3)\land\lozenge_{[0,60]}(x<24.0)$ not reach island and reach harbor
@@ V-B Synthetic Dataset
$p=(\alpha\leq x\leq\beta)\land(\gamma\leq y\leq\lambda),\$ (20)
@@
$\phi_{4}$ $=\lozenge_{[10,40]}(-5\leq x\leq-3\land-5\leq y\leq-3).$
Figure 3 shows some example trajectories in the synthetic dataset. The signal is required to arrive at some regions within the desired time interval specified by the STL specifications corresponding to the class. The synthetic dataset contains 5 classes labeled as $c_{1}$, $c_{2}$, $c_{3}$, $c_{4}$, $c_{5}$ with trajectories individually satisfy:
@@
fold $mSTL$ $dt\Delta$ $rnn$ $mSTL$ $dt\Delta$ $rnn$ $mSTL$ $dt\Delta$ $rnn$ $mSTL$ $dt\Delta$ $rnn$ $mSTL$ $dt\Delta$ $rnn$ $mSTL$ $dt\Delta$ $rnn$
1 0.00 0.00 0.00 64 324 67 0.00 0.00 0.00 63 1689 69 0.00 0.00 0.00 65 3236 67
2 0.00 0.00 0.00 63 299 68 0.00 0.00 0.00 62 1431 68 0.00 0.00 0.00 64 3220 67
3 0.00 0.00 0.00 65 427 68 0.00 0.00 0.00 62 1786 73 0.00 0.00 0.00 64 3360 66
4 0.00 0.00 0.00 63 395 66 0.00 0.00 0.00 63 1594 68 0.00 0.00 0.00 63 3821 67
5 0.00 0.00 0.00 63 359 67 0.00 0.00 0.00 63 1931 65 0.00 0.00 0.00 63 3353 67
0.00 0.00 0.00 64 361 67 0.00 0.00 0.00 63 1686 69 0.00 0.00 0.00 64 3398 67
@@
We conduct experiments on the datasets with sample sizes of $500$, $2500,$ and $5000$; these three datasets have $100$, $500,$ and $1000$ samples for each class, respectively. We also compare the performance with the decision-tree-based algorithm [23] and a recurrent neural network [24] on these datasets. The decision-tree-based method ($dt\Delta$) uses a decision tree with a maximum depth of 5. The recurrent neural network (denoted as $rnn$ in Table VI) has $4$ hidden layers, with the first layer as an LSTM layer.
@@
Table VI shows the results of these three methods on a 5-fold cross-validation. The initialization of the multi-class NN-TLI is consistent for each fold. The batch size and the number of training epochs are the same for neural networks, i.e., multi-class NN-TLI and $rnn$, across all folds. The ECOC matrix we use is the class-based coding matrix since $dt\Delta$ algorithm cannot implement the attribute learning. The results show that the multi-class NN-TLI, called $mSTL$ for short reference, can learn the STL formulae describing the properties of each class and achieve a low misclassification rate with high efficiency. The $mSTL$ and $rnn$ methods can achieve much higher efficiency than the $dt\Delta$, while $rnn$ cannot provide the characteristics of trajectories of each class. The decision tree-based algorithm $dt\Delta$ is able to learn the STL formulae. However, the training time of $dt\Delta$ increases with the sample size of the dataset. The reason for that is the computational bottleneck of the decision tree method, as it iterates all the input instances to find the optimal split at a given node. This step is repeated for each node of the tree [25]. Thus, the computational time grows with the total number of instances. However, the neural network method iterates over the selected batch data at each epoch, which is small relative to the total number of instances.
$f_{1}$ $f_{2}$ $f_{3}$ $f_{4}$
$c_{1}$ + + - -
$c_{2}$ + - + -
$c_{3}$ + + -
$c_{4}$ +
(a) The coding matrix $E_{obs}$
$f_{1}$ $f_{2}$ $f_{3}$ $f_{4}$
$c_{5}$ + - -
(b) The coding matrix $E_{pred}$
TABLE VII: The ECOC coding matrix for zero-shot learning.
To investigate the effectiveness of the margin in the loss function during the training process, we recorded the average absolute robustness value of the correctly classified trajectories shown in Fig 4. For comparison, we also trained a multi-class NN-TLI without margin in the loss function and recorded the average absolute robustness, also shown in Fig 4. It can be observed that the introduction of the margin to the loss can increase the absolute value of the average robustness. In other words, the margin can improve the separation between the trajectories during the training process and achieve better performance.
We also implement a zero-shot learning [26] to demonstrate the effectiveness of attribute-based recognition in our multi-class NN-TLI framework. We extract attributes from observed samples and employ them to predict the class of non-observed data through attribute composition. We generate $5$ classes of samples based on $4$ types of attributes, with their labels denoting possession of attributes, as shown in Table VII. The multi-class NN-TLI is trained over observed classes $c_{1}$, $c_{2}$, $c_{3}$, $c_{4}$ using the coding matrix $E_{obs}$ in Table VIIa and learn $4$ binary classifier $f_{1}$, $f_{2}$, $f_{3}$, $f_{4}$ correlating to $4$ kinds of attributes. These four attributes are interpreted by four STL formulae:
@@
$\psi_{4}$ $=\lozenge_{[1,20]}(-4.1<x<-0.2\land-3.8<y<-0.5),$
we then use these attributes to predict the unobserved class $c_{5}$ labeled as $E_{pred}$ in Table VIIb. We can achieve $0.00$ $MCR$ for classifying the trajectories in $c_{5}$. The multi-class NN-TLI can capture the attributes of the signals and interpret the characteristics through the learned STL formulae.
@@ VI Conclusions and Future Work
We propose a multi-class NN-TLI neural network designed for multi-class classification tasks on time-series data. The neural network can learn STL formulae to interpret signal attributes and classify signals based on their attribute representation. Our method is efficient, offering analysis from a general perspective and providing concise descriptions of the data. In the future, we will focus on determining the importance of attributes by learning the weight of each attribute.